A mainframe computer whose cost is k220,000 will depreciate to a scrap value of k12,000in 5 years.
(1)what is the book value of the computer at the end of the third year?
(2)how much more would the book value be at the end of the third year of the straight line method of depreciation had been used?

Answer by Boreal(15235)

exponential
P=Ae^kt
12000=220000e^5k
12/220=e^5k
ln both sides
-2.909=5k
k=-0.5817
-
after 3 years
P=220000*e^(-1.745)
P=k38,412.96 rounding at end
-
straight line
(0, 220)
(5, 12)
slope is 208/-5=-41.6 thousands
It decreases k41,600 a year for 3 years or 124,800 less.
220,000-124,800=K95,200
The book value would be k56,787 more